How fast must an object be going if its rest length is 3.4 times as long as its observed length?
Answer #1

The formula for observed length is given by:

Observed length = (Length at rest)*{1-(v/c)2}1/2

Now, Length at rest = 3.4*Observed Length

Therefore, 1 = 3.4*{1-(v/c)2}1/2

1 = 11.56*{1-(v/c)2}

{1-(v/c)2} = 1/11.56 = 0.0865

(v/c)2 = 0.9135

v/c = 0.956

v = 0.956*c = 0.956 * 3 * 108 m/s

v = 2.87*108 m/s

Therefore, the object must be moving at 2.87*108 m/s.

I hope i was able to answer your question. If you need anymore explanation, please ask in comments. I would be happy to help.

Please rate the solution if you are satisfied with the solution.